OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (5, 6) and (9, 14)?

OpenStudy (dtan5457):

The formula for slope intercept is y=mx+b You will need to get the slope first... which is \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

which is 9-5=4 and 14-6=8

OpenStudy (dtan5457):

correct, now divide

OpenStudy (anonymous):

what do i divide?

OpenStudy (dtan5457):

8/4, just look at the slope formula, y/x

OpenStudy (dtan5457):

Any confusion ?

OpenStudy (anonymous):

so its 2?

OpenStudy (dtan5457):

Yeah, the slope is 2. Now you have to solve for "b" in the y=mx+b Plug in a point to the formula let's say (5,6) (in your question) 6=2(5)+b solve for b like regular algebra

OpenStudy (anonymous):

so b would equal 10 or no

OpenStudy (dtan5457):

6=10+b subtract the 10 on both sides

OpenStudy (anonymous):

so it = 4